# Robots Can Multitask Too: Integrating a Memory Architecture and LLMs for Enhanced Cross-Task Robot Action Generation

### v1

@@ I INTRODUCTION
Despite the physical limitations due to their embodiment, humanoid robots are particularly effective tools because of their anthropomorphic shape, which can significantly improve the versatility and effectiveness of robots, especially in environments designed for human interaction [1]. Moreover, the humanoid physical shape supports collaborating with humans whose legibility and predictability of robot actions are fostered by the robot’s physicality, with positive impacts on the safety and trust during the interaction [2]. Inspired by the effectiveness and adaptability of human cognitive abilities, cognitive robotics aims to draw from more than embodiment when designing robotic platforms. Key attributes like human perception, selective attention, different types of memory, and other facets of human cognition are essential to improve the autonomy, flexibility, and capabilities of robots [3].
@@ II Related Work
Integrating sensory input with memory is essential in artificial systems as it enables them to retain sensory information for environment recognition and adaptive responses [11]. Moreover, memory plays a vital role in humanoid robots since it brings together conceptual knowledge gained through experience with the robot’s behavior, enabling adjusting to complex interactive scenarios and tasks [12]. Memory-based architectures have various applications such as transferring robot manipulation skills across different human environments [13], memory-powered incremental learning with human-in-the-loop corrections [14], and supporting tasks like action-planning for localizing a target and navigating towards it [15]. Existing work incorporates models of memory into robot tasks like planning with occluded objects [16], reasoning about object permanence [17], and learning human personal attributes like voices and faces [18]. However, such approaches model memory representations and reasoning as distinct components or separate processes while embedding memory directly within a reasoning backbone, like an LLM, is beneficial for context-aware decision-making.
@@
Figure 2: A simplified overview of our system’s workflow. The workflow starts when the NICOL robot receives a given task.
@@ III-A Robotic Platform: NICOL
Our system runs on the Neuro-Inspired COLlaborator (NICOL) [27] built in the Knowledge Technology group at the University of Hamburg. NICOL is a tabletop robotic platform with a humanoid head and two 8 DoF robotic arms. The robot features social cues like facial expressions as well as object and arm manipulation capabilities like grasping and neuro-genetic inverse kinematics [28]. The NICOL platform is based on the Robot Operating System (ROS), facilitating integration with various sensory modules and deep learning models. It also supports open communication and collaboration with human users through an LLM-powered modular architecture [29] that combines sensory input like visually grounded open-vocabulary object detection (ViLD [30]) and speech recognition (Whisper [31]) with actions like arm and head manipulation. Thus, it renders the robot with social and cognitive skills through its grounded LLM. The NICOL platform can also parse and translate text-defined actions such as “point to the banana” into corresponding physical actions performed by the robot’s arm, i.e., an extended arm and extended index finger gesture to indicate the location of the banana on the table. In our approach, we extend NICOL’s skills, however, in contrast to previous work, we focus on adaptive action generation in robot multi-task scenarios.
@@ III-B NICOL’s Visual Perception and Action Parsing
We use NICOL’s 4K camera located inside the robot’s left eye to capture the environment visually and the objects on the table. We use the ViLD [30] object detector, enabling zero-shot object detection and recognition without pre-training. The NICOL’s ViLD object detector has been optimized to work in real-time on continuous camera streams with high efficiency and reliability [29]. The detector generates bounding boxes for the objects on the table and their locations. All objects detected outside the area of interest, i.e., the robot’s table, are removed. The objects detected are expressed by text so all object labels can be streamlined within the system’s pipeline, making it compatible with an LLM application. NICOL can also parse textual LLM-generated actions and translate them into robot physical actions [29]. The robot can perform arm manipulations like pointing to objects on the table, grasping objects, and pushing objects to the user. The robot also has other head manipulation and facial expression capabilities. However, our system utilizes only the robot’s arm manipulation skills in its action and task specifications.
@@ III-C Proposed Architecture
Prompt 1: Task specifications as given to the coordinator LLM in the base prompt. Sorting task: Your task is to sort objects in different boxes based on their properties. Fruits should go to box 1 while kitchenware and containers should go to box 2. For this task, you can place objects in the corresponding boxes using appropriate actions from the list of functions, for example: $<$move_to_box_1(banana)$>$.
@@
Prompt 3: Memory functions for triggering the worker LLM. $<$retrieve_working_memory(task)$>$: Given a string of a task name, use the following action format $<$retrieve_working_memory(task_name)$>$ to retrieve the working memory. For instance, if you want to retrieve the working memory for sorting, use $<$retrieve_working_memory(sorting)$>$. Ensure that the output follows this format exactly.
retrieve_declarative_memory(task)$>$: Given a string of a task name, use the following action format $<$retrieve_declarative_memory(task_name)$>$ to retrieve the declarative memory. For instance, if you want to retrieve the declarative memory for sorting, use $<$retrieve_declarative_memory(sorting)$>$. Ensure that the output follows this format exactly.
@@ III-D LLMs for Instructions and Reasoning
III-E Working Memory
The working memory also includes a selective object state extracted by combining the information of the declarative and procedural memory. Specifically, it receives all the objects on the table and outputs only the task-relevant objects, representing the robot’s visual selective attention. In Prompt III-E, we highlight an example of the sorting task (fruits and dishware items must be placed in different bins). The first part of the prompt is the task description, which is then combined with the visual input identifying the objects: apple, banana, cup, bowl, baseball, pear. The prompt’s final part sets rules for the worker LLM on output formulation, ensuring the working memory remains consistent before passing it on back to the coordinator LLM. As a result, only necessary objects are retained, i.e., objects not task-relevant are removed from the working memory (the baseball in the given example).
Prompt 4: An example prompt to extract the working memory (sorting task). Prompt Input:
Sorting Task: the task is to move fruits to box 1, containers and kitchenware objects to box 2. $\longrightarrow$ Task Description
@@ III-E Working Memory
Prompt 5: An example prompt to extract the declarative memory (sorting task). Prompt Input:
@@ III-F Declarative Memory
When the robot modifies the environment with object manipulation actions, these changes are logged into the system’s declarative memory, ensuring that the robot’s internal state through its memory is precisely synchronous with the external world. Since all the robot actions are recorded, the system always has knowledge of the interaction history over any duration of time. We store this memory persistently as log files, one file per task. Similar to working memory, declarative memory is specialized to one task and does not contain information about other tasks. As a task newly starts, the interaction history is empty and data is accumulated as the task progresses. If the task gets interrupted, i.e., the robot is asked to switch to another one, the interaction history is kept safe. When the robot continues the interrupted task, it retrieves essential information like previous actions, environment state, and task state at the interruption point.
The worker LLM is responsible for creating and maintaining the logs of the declarative memory. An example using the sorting task is shown in Prompt III-E (for simplicity, we only show the information extraction of box 1). The worker LLM is given the persistent task logs and prompted to extract the remaining objects on the table (environment state) and the objects in box 1 and box 2 (task state). Similar to the working memory, the prompt has a clearly defined output structure for consistency of the information in the system across the two LLMs. The log file is structured by entries ordered by the executed actions, which allows the LLM to efficiently identify the most recent environment and task states.
@@ IV Experiments and Evaluation
We evaluate our system using five robotic tasks, each requiring a set of objects as input and an appropriate sequence of robot actions as output to achieve the task expected goal. Each task involves both a cognitive aspect, where the robot must analyze the objects and reason about the task definition to identify the ones required to achieve it, and a physical aspect, in which the robot performs the appropriate object manipulation function. The objects necessary to perform tasks vary per task, however, we limit the total input to six objects since our early experiments showed that an input size of six objects is most reasonable given the limited number of tokens the LLM can process within its context window. We use YCB objects for the tasks in addition to colored cubes for the tower task. An overview of the tasks is as follows: 1) Sorting Task: the robot places fruits and dishwasher items in different containers, 2) Arrangement Task: the fruits must be arranged by placing them inside a bowl, 3) Pointing Task: the robot points to all the yellow objects on the table and then to all the red objects, 4) Recipe Task: objects that are required to make a jello recipe with banana topping are given to the user, and 5) Tower Task: the robot builds a tower by stacking only colored cubes on top of each other. A summary of the tasks is provided in Table I. Our system runs on the NICOL robot, but for evaluation, we disable the robot’s physical control parts of our system since we focus on evaluating its action-generating aspect and not the robot’s arm manipulation.
@@ IV-C Results for Consecutive Tasks
In this setup, the system has knowledge of all the defined tasks and possible actions through its declarative memory. Hence, the LLM must not only reason about the task requirements but also choose the appropriate actions. In this experiment, we observe that running the tasks consequently presents a considerable challenge to the LLMs when tuned without the working memory (cf. Table III). The system achieves average performance for some of the tasks (sorting and recipe), however, the success rate and retention drop significantly for most tasks compared to the standalone experiment. The pointing task is particularly challenging since the LLM must reason about the order of objects, not only their type, as well as the appropriate pointing action. Similarly, the tower task is challenging to the system, and we observe failures in correctly reasoning that black and white cubes (not colored) must not be stacked in the tower.
In contrast, we report a noticeable boost in the success rate and retention metrics when running the experiment with the working memory enabled (see Table III). Both LLM variants perform comparably and even reach a perfect score for some of the tasks, e.g., arrangement. Noticeably, Llama 3 outperforms ChatGPT 3.5 on the sorting task due to errors in the coordinator bypassing the worker LLM output and generating redundant unnecessary actions, negatively influencing its overall success and retention values. Overall, both LLMs achieve success rates that approximate the baseline (standalone setup) and even surpass it for tasks like arrangement and recipe (see Fig. 5). This experiment shows a significant improvement in performance due to the utilization of the worker LLM and working memory, thus, suggesting that the coordinator LLM becomes a bottleneck in action generation over a longer period of interaction time.
@@ IV-D Results for Intervened Tasks
We evaluate the declarative and working memory by creating intervention points, i.e., interruptions, in which the execution of a task is paused, and the system must switch to the next task. Each trial in this experiment involves two execution rounds. In the first round, only 50% progress of each task is accomplished, and in the second, the system resumes every paused task from the interruption points until full completion. For consistency, we run the tasks by their definition order: sorting, arrangement, pointing, recipe, and tower. However, the task order does not influence the experiment since each task has its own input and runs independently, i.e., one task success or failure does not impact other tasks’ performance. We run the intervention mode twice: without memory and with declarative and working memory. Similar to the consecutive mode, the tasks are challenging to both LLMs without memory (cf. Table IV). Noticeably, the recipe task performance drops for GPT 3.5 due to errors in the coordinator LLM, generating noisy actions on wrong objects. Moreover, the performance of Llama 3 is below average for all tasks. The main error source is the LLM “forgetting” the task specifications due to the longer interaction duration, and thus, generating false actions or no actions. Also, the logical task reasoning was challenging, and the coordinator LLM generated incorrect actions, e.g., pointing to an object neither yellow nor red, placing a non-fruit object in the bowl, or giving the wrong object in the recipe task.
@@
Here again, our experiment demonstrates noticeable improvement in the success rate, task retention, and environment retention scores across both LLMs (see Table IV). The scores for the pointing, recipe, and tower tasks are comparable for both models and approximate the baseline (see Fig. 6). Interestingly, Llama 3 scored higher success and retention values in the sorting task due to the task grounding of GPT 3.5 leading to faulty commonsense reasoning, thus, occasionally generating wrong actions such as $<$move_to_box_2(pear)$>$ instead of $<$move_to_box_1(pear)$>$. We also observed Lama 3 generating a wrong list of objects in the working memory of the arrangement task, which explains the slight drop in its performance. However, when comparing against the previous experiment (consecutive with memory), this seems to be a challenge only when the system must provide this information in the second round of task execution. We hypothesize this to be due to the context size of Llama 3 being roughly half the size of this ChatGPT model (8k vs 16k) since this mode requires multiple rounds. We also observed high variability in the output with these answers, hinting that the parameters chosen for Llama 3 could be further optimized for more stable output (ex: we used $top\_p$ of 1 for both models).
@@ V Discussion and Conclusion
Our experiments showed an inherent challenge for the current generation of LLMs (ChatGPT and Llama) to manage different tasks and goals at the same time by relying on the LLM’s own internal memory or the context of previous dialogues in the prompt. However, our implemented memory process (declarative, procedural, and working memory) demonstrated a significant boost in the performance of both models compared to our baseline. In future work, memory-based architectures like our proposed one can be expected to be beneficial for human-robot collaboration besides action generation. For example, we plan research on managing high-level goals in multi-party robotic scenarios, where humans are interacting with the robot, each having their own defined goal. As LLMs with larger context windows become more available and accessible, we will also evaluate our system with a larger number of tasks and input objects.

### v2

@@ I INTRODUCTION
Despite the physical limitations due to their embodiment, humanoid robots are particularly effective tools because of their anthropomorphic shape, which can significantly improve the versatility and effectiveness of robots, especially in environments designed for human interaction [1]. Moreover, the humanoid physical shape of the robot supports collaboration with humans by making its actions more legible and predictable, which positively impacts safety and trust during interactions [2]. Inspired by the effectiveness and adaptability of human cognitive abilities, cognitive robotics aims to draw from more than embodiment when designing robotic platforms. Key attributes like human perception, selective attention, different types of memory, and other facets of human cognition are essential to improve the autonomy, flexibility, and capabilities of robots [3].
@@ II Related Work
Integrating sensory input with memory is essential in artificial systems as it enables them to retain sensory information for environment recognition and adaptive responses [11]. Moreover, memory plays a vital role in humanoid robots since it brings together conceptual knowledge gained through experience with the robot’s behavior, enabling adjustment to complex interactive scenarios and tasks [12]. Memory-based architectures have various applications such as transferring robot manipulation skills across different human environments [13], memory-powered incremental learning with human-in-the-loop corrections [14], and supporting tasks like action-planning for localizing a target and navigating towards it [15]. Existing work incorporates models of memory into robot tasks like planning with occluded objects [16], reasoning about object permanence [17], and learning human personal attributes like voices and faces [18]. However, such approaches model memory representations and reasoning as distinct components or separate processes while embedding memory directly within a reasoning backbone, like an LLM, is beneficial for context-aware decision-making.
@@
Figure 2: A simplified overview of our system’s workflow. The workflow starts when the NICOL robot receives a given task. The robot’s sensory and task inputs are then fed to our proposed LLM-powered architecture.
@@ III-A Robotic Platform: NICOL
Our system runs on the Neuro-Inspired COLlaborator (NICOL) [27] built in the Knowledge Technology group at the University of Hamburg. NICOL is a tabletop robotic platform with a humanoid head and two 8 DoF robotic arms. The robot features social cues like facial expressions as well as object and arm manipulation capabilities like grasping and neuro-genetic inverse kinematics [28]. The NICOL platform is based on the Robot Operating System (ROS), facilitating integration with various sensory modules and deep learning models. It also supports open communication and collaboration with human users through an LLM-powered modular architecture [29] that combines sensory input like visually grounded open-vocabulary object detection (ViLD [30]) and speech recognition (Whisper [31]) with actions like arm and head manipulation. Thus, it equips the robot with social and cognitive skills via its grounded LLM. In our approach, we extend NICOL’s skills, but unlike previous work, we focus on adaptive action generation in robot multi-task scenarios.
@@ III-B NICOL’s Visual Perception and Action Parsing
We use NICOL’s 4K camera located inside the robot’s left eye to capture the environment visually and the objects on the table. We use the ViLD [30] object detector, enabling zero-shot object detection and recognition without pre-training. The NICOL’s ViLD object detector has been optimized to work in real-time on continuous camera streams with high efficiency and reliability [29]. The detector generates bounding boxes for the objects on the table and their locations. All objects detected outside the area of interest, i.e., the robot’s table, are removed. The objects detected are expressed by text so all object labels can be streamlined within the system’s pipeline, making it compatible with an LLM application. NICOL can also parse textual LLM-generated actions and translate them into robot physical actions [29]. The robot can perform arm manipulations like pointing to objects on the table, grasping objects, and pushing objects to the user. For example, “point to the banana” can be parsed and translated into corresponding physical actions performed by the robot’s arm, i.e., an extended arm and extended index finger gesture to indicate the location of the banana on the table. The robot also has other head manipulation and facial expression capabilities. However, our system utilizes only the robot’s arm manipulation skills in its action and task specifications.
@@ III-C Proposed Architecture
Prompt 1: Task specifications as given to the coordinator LLM in the base prompt. Separating task: Your task is to sort objects in different boxes based on their properties. Fruits should go to box 1 while kitchenware and containers should go to box 2. For this task, you can place objects in the corresponding boxes using appropriate actions from the list of functions, for example: $<$move_to_box_1(banana)$>$.
@@
Prompt 3: Memory functions for triggering the worker LLM. $<$retrieve_working_memory(task)$>$: Given a string of a task name, use the following action format $<$retrieve_working_memory(task_name)$>$ to retrieve the working memory. For instance, if you want to retrieve the working memory for separating, use $<$retrieve_working_memory(separating)$>$. Ensure that the output follows this format exactly.
retrieve_declarative_memory(task)$>$: Given a string of a task name, use the following action format $<$retrieve_declarative_memory(task_name)$>$ to retrieve the declarative memory. For instance, if you want to retrieve the declarative memory for separating, use $<$retrieve_declarative_memory(separating)$>$. Ensure that the output follows this format exactly.
@@ III-D LLMs for Instructions and Reasoning
Prompt 4: An example prompt to extract the working memory (separating task). Prompt Input:
Separating Task: the task is to move fruits to box 1, containers and kitchenware objects to box 2. $\longrightarrow$ Task Description
@@
Prompt 5: An example prompt to extract the declarative memory (separating task). Prompt Input:
@@
III-E Working Memory
The working memory also includes a selective object state extracted by combining the information of the declarative and procedural memory. Specifically, it receives all the objects on the table and outputs only the task-relevant ones, representing the robot’s visual selective attention. In Prompt III-D, we highlight an example of the separating task (fruits and dishware items must be placed in different bins). The first part of the prompt is the task description, which is then combined with the visual input identifying the objects: apple, banana, cup, bowl, baseball, pear. The prompt’s final part sets rules for the worker LLM on output formulation, ensuring the working memory remains consistent before passing it back to the coordinator LLM. As a result, only necessary objects are retained, i.e., objects not task-relevant are removed from the working memory (the baseball in the given example).
@@ III-F Declarative Memory
When the robot modifies the environment with object manipulation actions, these changes are logged into the system’s declarative memory, ensuring that the robot’s internal state through its memory is precisely synchronous with the external world. Since all the robot actions are recorded, the system always has knowledge of the interaction history over any duration of time. We store this memory persistently as log files, one file per task. Similar to working memory, declarative memory is specialized to one task and does not contain information about other tasks. As a task newly starts, the interaction history is empty and data is accumulated as the task progresses. Although the memory is extended dynamically, since the robot generates one action per object, the memory size (number of entries) can be at most equal to the number of objects. If the task gets interrupted, i.e., the robot is asked to switch to another one, the interaction history is kept safe. When the robot continues the interrupted task, it retrieves essential information like previous actions, environment state, and task state at the interruption point.
The worker LLM is responsible for creating and maintaining the logs of the declarative memory. An example using the separating task is shown in Prompt III-D (for simplicity, we only show the information extraction of box 1). The worker LLM is given the persistent task logs and prompted to extract the remaining objects on the table (environment state) and the objects in box 1 and box 2 (task state). Similar to the working memory, the prompt has a clearly defined output structure for consistency of the information in the system across the two LLMs. The log file is structured by entries ordered by the executed actions, which allows the LLM to efficiently identify the most recent environment and task states.
@@ IV Experiments and Evaluation
We evaluate our system using five robotic tasks, each requiring a set of objects as input and an appropriate sequence of robot actions as output to achieve the task expected goal. Each task involves both a cognitive aspect, where the robot must analyze the objects and reason about the task definition to identify the ones required to achieve it, and a physical aspect, in which the robot performs the appropriate object manipulation function. The objects necessary to perform tasks vary per task, however, we limit the total input to six objects since our early experiments showed that an input size of six objects is most reasonable given the limited number of tokens the LLM can process within its context window. We use YCB objects for the tasks in addition to colored cubes for the tower task. An overview of the tasks is as follows: 1) Separating Task: the robot places fruits and dishwasher items in different containers, 2) Arrangement Task: the fruits must be arranged by placing them inside a bowl, 3) Pointing Task: the robot points to all the yellow objects on the table and then to all the red objects, 4) Recipe Task: objects for making a jello recipe with banana topping are given to the user, and 5) Tower Task: the robot builds a tower by stacking only colored cubes on top of each other. A summary of the tasks is provided in Table I. Our system runs on the NICOL robot, but for evaluation, we disable the robot’s physical control since we focus on evaluating our system’s action-generating aspect and not the NICOL platform components.
@@ IV-C Results for Consecutive Tasks
In this setup, the system has knowledge of all the defined tasks and possible actions through its declarative memory. Hence, the LLM must not only reason about the task requirements but also choose the appropriate actions. In this experiment, we observe that running the tasks consequently presents a considerable challenge to the LLMs when tuned without the working memory (cf. Table III). The system achieves average performance for some of the tasks (separating and recipe), however, the success rate and retention drop significantly for most tasks compared to the standalone experiment. The pointing task is particularly challenging since the LLM must reason about the order of objects, not only their type, as well as the appropriate pointing action. Similarly, the tower task is challenging to the system, and we observe failures in correctly reasoning that black and white cubes (not colored) must not be stacked in the tower.
In contrast, we report a noticeable boost in the success rate and retention metrics when running the experiment with the working memory enabled (see Table III). Both LLM variants perform comparably and even reach a perfect score for some of the tasks, e.g., arrangement. Noticeably, Llama 3 outperforms ChatGPT 3.5 on the separating task due to errors in the coordinator bypassing the worker LLM output and generating redundant unnecessary actions, negatively influencing its overall success and retention values. Overall, both LLMs achieve success rates that approximate the baseline (standalone setup) and even surpass it for tasks like arrangement and recipe (see Fig. 5). This experiment shows a significant improvement in performance due to the utilization of the worker LLM and working memory, thus, suggesting that the coordinator LLM becomes a bottleneck in action generation over a longer period of interaction time.
@@ IV-D Results for Intervened Tasks
We evaluate the declarative and working memory by creating intervention points, i.e., interruptions, in which the execution of a task is paused, and the system must switch to the next task. Each trial in this experiment involves two execution rounds. In the first round, only 50% progress of each task is accomplished, and in the second, the system resumes every paused task from the interruption points until full completion. The 50% intervention is chosen for simplicity; however, the system behaves similarly at any interruption point. For consistency, we run the tasks by their definition order: separating, arrangement, pointing, recipe, and tower. However, the task order does not influence the experiment since each task has its own input and runs independently, i.e., one task success or failure does not impact other tasks’ performance. We run the intervention mode twice: without memory and with declarative and working memory. Similar to the consecutive mode, the tasks are challenging to both LLMs without memory (cf. Table IV). Noticeably, the recipe task performance drops for GPT 3.5 due to errors in the coordinator LLM, generating noisy actions on wrong objects. Moreover, the performance of Llama 3 is below average for all tasks. The main error source is the LLM “forgetting” the task specifications due to the longer interaction duration, and thus, generating false actions or no actions. Also, the logical task reasoning was challenging, and the coordinator LLM generated incorrect actions, e.g., pointing to an object neither yellow nor red, placing a non-fruit object in the bowl, or giving the wrong object in the recipe task.
@@
Here again, our experiment demonstrates noticeable improvement in the success rate, task retention, and environment retention scores across both LLMs (see Table IV). The scores for the pointing, recipe, and tower tasks are comparable for both models and approximate the baseline (see Fig. 6). Interestingly, Llama 3 scored higher success and retention values in the separating task due to the task grounding of GPT 3.5 leading to faulty commonsense reasoning, thus, occasionally generating wrong actions such as $<$move_to_box_2(pear)$>$ instead of $<$move_to_box_1(pear)$>$. We also observed Lama 3 generating a wrong list of objects in the working memory of the arrangement task, which explains the slight drop in its performance. However, when comparing against the previous experiment (consecutive with memory), this seems to be a challenge only when the system must provide this information in the second round of task execution. We hypothesize this to be due to the context size of Llama 3 being roughly half the size of this ChatGPT model (8k vs 16k) since this mode requires multiple rounds. We also observed high variability in the output with these answers, hinting that the parameters chosen for Llama 3 could be further optimized for more stable output (ex: we used $top\_p$ of 1 for both models).
@@ V Discussion and Conclusion
Our experiments showed a challenge for the current LLMs (ChatGPT and Llama) to manage different tasks and goals at the same time by relying on the LLM’s own internal memory or the context of previous dialogues in the prompt. However, our implemented memory process (declarative, procedural, and working memory) demonstrated a significant boost in the performance of both models compared to our baseline. Although LLMs enable the flexibility of task definition and reasoning without explicit programming, their fixed context size can restrict the number of tasks and objects that can be managed simultaneously. Also, our experiments showed that LLMs are still prone to errors like occasionally generating redundant or incorrect actions. However, in future work, memory-based architectures like our proposed one can be expected to be beneficial for human-robot collaboration besides action generation. For example, we plan research on managing high-level goals in multi-party robotic scenarios, where humans are interacting with the robot, each having their own goal. As LLMs with larger context windows become more available and accessible, we will also evaluate our system with a larger number of tasks and input objects.